OpenStudy (anonymous):

Please Help! MEDAL WILL BE REWARDED! lim sec((pi(x))/(6)) x->3+

OpenStudy (anonymous):

sec (πx/6) = 1/cos(πx/6) =1/cos(π/2) =1/0+=+∞